OpenStudy (anonymous):

Prove: sin θ - sin θ•cos2 θ = sin3 θ. You must show all work.

OpenStudy (anonymous):

idk even where to begin

myininaya (myininaya):

Just thought Maybe you can take the right hand side and show the left hand side Try to play with this thought: sin(3x) is sin(2x+x)

myininaya (myininaya):

That should be a hint to use the sum identity for sin

OpenStudy (idealist10):

Memorize the identities for cos 2*theta

myininaya (myininaya):

Though I could be wrong but I don't think this is an identity.

myininaya (myininaya):

One way to find out is to find a counterexample

myininaya (myininaya):

try theta=pi/2 does not work

OpenStudy (idealist10):

sin(2x)=2*sin(x)*cos(x)

OpenStudy (idealist10):

cos(2x)=2cos^2 (x)-1

myininaya (myininaya):

Maybe the problem has been written incorrectly.

myininaya (myininaya):

Unless the instructions say to: Prove if true and give a counterexample if false. If so then I have accidentally given you the answer.

OpenStudy (idealist10):

cos(2x)=1-2sin^2 (x) so what's sin(x)*cos(2x)?